OpenStudy (anonymous):

In parallelogram PQRS, find the value of angle SPR if angle RQP is 75° and angle PRS is 39°

OpenStudy (katrinakaif):

Can you draw it?

OpenStudy (anonymous):

yea give me a sec

OpenStudy (saifoo.khan):

I have solved it. xD

OpenStudy (anonymous):

really???

OpenStudy (saifoo.khan):

Ye!

OpenStudy (saifoo.khan):

yea*

OpenStudy (saifoo.khan):

let me draw it.

OpenStudy (saifoo.khan):

can u solve it now?

OpenStudy (anonymous):

75?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

x + 39 + 75 = 180 Solve for x.. :)

OpenStudy (anonymous):

umm ok x + 114=180 -114 -114 x = 66?

OpenStudy (saifoo.khan):

Yes!!!!

OpenStudy (anonymous):

yaaaaaaaaaaaaaay!!!!!!!!!! lozl

OpenStudy (saifoo.khan):

LOL.